# A Statistical Learning Approach for Feature-Aware Task-to-Core Allocation in Heterogeneous Platforms

### v1

@@ Abstract
Optimizing task-to-core allocation can substantially reduce power consumption in multi-core platforms without degrading user experience. However, many existing approaches overlook critical factors such as parallelism, compute intensity, and heterogeneous core types. In this paper, we introduce a statistical learning approach for feature selection that identifies the most influential features—such as core type, speed, temperature, and application-level parallelism or memory intensity—for accurate environment modeling and efficient energy optimization. Our experiments, conducted with state-of-the-art Linux governors and thermal modeling techniques, show that correlation-aware task-to-core allocation lowers energy consumption by up to 10% and reduces core temperature by up to 5C compared to random core selection. Furthermore, our compressed, bootstrapped regression model improves thermal prediction accuracy by 6% while cutting model parameters by 16%, yielding an overall mean square error reduction of 61.6% relative to existing approaches. We provided results based on superscalar Intel Core i7 12th Gen processors with 14 cores, but validated our method across a diverse set of hardware platforms and effectively balanced performance, power, and thermal demands through statistical feature evaluation.
Keywords Energy Optimization, Heterogeneous Platforms, Statistical Feature Evaluation, Task-to-Core Allocation
@@ 1 Introduction
Dynamic voltage and frequency scaling (DVFS) and task-to-core allocation are critical techniques for optimizing the performance and energy efficiency of embedded systems. Since power consumption is exponentially related to voltage and frequency, an effective DVFS strategy can reduce power consumption by up to 75% without impacting user experience Ratković et al. (2015). However, different workloads exhibit varying thermal and power consumption behaviors depending on task-to-core allocation strategies. In multiprocessor systems, cores with diverse characteristics, such as high-performance cores, low-power cores, and GPUs, offer unique performance trade-offs. Suboptimal core allocation can lead to thermal throttling, reduced reliability, and shorter device lifespan Hosseinimotlagh and Kim (2019). Furthermore, the cooling costs associated with overheating can account for approximately \$3 per watt of heat dissipation et al. (2003).
To address these challenges, existing approaches rely primarily on historical workload performance and sensor data to guide future processor behavior Maity et al. (2022). However, these approaches often fail to generalize in varying workload conditions and hardware architectures.
Multi-Core heterogeneous processors, which are widely used in automotive, telecommunications, and consumer electronics, must operate within strict power and thermal constraints while maintaining high performance. In heterogeneous architectures, DVFS and task-to-core allocation are essential for managing these constraints by optimizing the trade-off between power efficiency and performance. Feature selection plays a crucial role in identifying key system parameters, such as frequency, temperature, and voltage, which influence thermal stability and energy efficiency Shekarisaz et al. (2021). For example, in real-time automotive systems, DVFS strategies can minimize energy consumption without sacrificing response times, whereas in mobile devices, efficient task-to-core allocation extends battery life by utilizing low-power cores for background tasks.
Despite these advancements, the reliance on real-world data collection for policy optimization introduces computational overhead and inaccuracies due to hardware sampling delays. A promising alternative is the use of statistical learning models to infer system behavior from collected data, reducing reliance on extensive real-world testing. Properly trained models can capture the stochastic nature of embedded environments and provide accurate control signal generation Liu et al. (2021). This work proposes a statistical learning approach to develop efficient models using data augmentation and feature selection, thus enhancing energy-efficient task scheduling.
To the best of our knowledge, limited work has been done in applying statistical feature selection for DVFS policy optimization. Previous efforts, such as the use of extreme value theorem (EVT) to estimate upper bounds of energy consumption and execution time Cazorla et al. (2019); Davis and Cucu-Grosjean (2019); Reghenzani et al. (2020), lack a comprehensive analysis of feature importance. Approaches such as XGBoost and decision trees have been employed to evaluate application-specific latency and energy features Liu et al. (2021); Sasaki et al. (2007), but a more robust, lightweight model suitable for multi-core platforms is needed.
This paper contributes the following:
@@
    A novel evaluation of feature importance for environment modeling and compressed learning for task-to-core allocation.
@@
    A correlation-aware task-to-core allocation approach to reduce temperature and energy consumption by strategically assigning tasks to uncorrelated cores.
@@
    A data augmentation strategy using bootstrapping to enhance model accuracy while reducing sample collection overhead.
@@
    Empirical validation against state-of-the-art thermal modeling methods, demonstrating up to 10% energy savings and $5^{\circ}$ temperature reduction. Our compressed regression model reduces prediction error by 6% while cutting the number of parameters by 16%, achieving a 61.6% reduction in mean square error.
@@ 2 Motivation and Challenges
Feature selection is a crucial data processing strategy in statistical learning, addressing challenges such as the curse of dimensionality, large data management, and performance unpredictability in multi-core platforms.
@@ 2.1 Curse of Dimensionality
Feature selection plays a vital role in mitigating the curse of dimensionality. As the number of features increases, the data space grows exponentially, leading to sparsity in high-dimensional spaces Li et al. (2017). This phenomenon can result in model overfitting, which degrades performance when predicting unseen data. In the context of heterogeneous systems, the increasing number and variety of processor cores—comprising performance, low-energy, and GPU cores, each characterized by frequency, temperature, and performance features—contribute to the expansion of the state space. As a result, efficient feature selection is necessary to reduce complexity and improve generalization.
@@ 2.2 Large Data Management
Feature selection is essential for handling the growing volume of data in DVFS and task-to-core allocation strategies. These strategies may rely on either static or streaming data, where static data consists of fixed historical samples, while streaming data continuously adapts to new input. With the increasing prevalence of streaming data, memory management becomes critical, as the unpredictable volume of incoming data can overwhelm system resources. Retaining unnecessary features significantly increases data storage and processing overhead, making feature selection a key factor in optimizing processor models and policy-learning algorithms.
@@ 2.3 Performance Unpredictability
Energy consumption in multi-core processors depends on multiple features, including frequency, voltage, temperature, and performance data from each core. As shown in Figure 1, energy consumption can vary by an order of magnitude when the same frequency and number of cores are used across different processors. This variability highlights the importance of feature selection in identifying the most significant predictors of energy consumption across different platforms, frequencies, and core configurations. Understanding which features retain their relevance across diverse conditions aids in developing a global policy for DVFS and task-to-core allocation.
Figure 1: Energy consumption variation across three different processors with identical frequency and core count settings: Intel Core i7 8th gen. (Corei78) with 4 cores, Intel Core i7 8th gen. (Corei712) with 14 cores, and Intel Xeon 2680 v3 (Xeon) with 12 cores. The results are shown for three different OpenMP benchmarks.
@@ 3 Design Methodology
Our objective is to develop a multi-stage methodology that selects the most critical features and applies these insights to optimize thermal behavior and energy efficiency in multi-core processors. In what follows, we present three feature selection approaches—filter-based, wrapper-based, and embedded—and show how their combined use informs an intelligent task-to-core allocation algorithm.
3.1 Filter-Based Feature Selection
Filter methods assess the relevance of features by examining intrinsic properties of the data without involving any learning algorithms. One common technique in filter methods is to evaluate the correlation between features and the target variable or among the features themselves. In our study, we utilize the Pearson correlation coefficient Sedgwick (2012) to quantify the linear relationship between core temperatures, which can indicate adjacency and potential heat transfer between cores.
The value of $r_{ij}$ ranges from $-1$ to $1$, where $1$ indicates a perfect positive linear correlation, $-1$ indicates a perfect negative linear correlation, and 00 signifies no linear correlation between the temperatures of cores $i$ and $j$. A high positive correlation suggests that the temperatures of the two cores rise and fall together, possibly due to physical proximity and shared thermal characteristics.
Correlation-Aware Task-to-Core Allocation Algorithm.
  1. 1.
  2. 2.
  3. 3.
  4. 4.
  5. 5.
This algorithm dynamically adapts to the thermal behavior of the system, effectively distributing the thermal load more evenly across the cores. Figure 3 presents a comparison of average temperature and energy consumption between correlation-based and random core selection across different governors. While random core allocation is generally expected to yield better results due to its unbiased distribution, the correlation-based approach demonstrates comparable performance, even when allocating a small number of cores to each workload. In scenarios where the order of core allocation is crucial—such as selecting 10 out of 14 available cores—the correlation-aware allocation method proves to be more effective in reducing both energy consumption and temperature, offering an advantage over random allocation.
3.2 Wrapper-Based Feature Selection
Wrapper methods evaluate subsets of features using a predictive model. Because they account for feature interactions, they typically yield higher accuracy than filter methods in finding the importance of the features on a specific feature parameter but may be more computationally expensive.
We employ the backward stepwise selection algorithm, which starts with all available features and iteratively removes the least significant feature based on a specified criterion. In our case, we use the Ordinary Least Squares (OLS) regression model to predict the target variables (energy consumption and average temperature) and assess the significance of the characteristics using statistical tests.
Backward Stepwise Selection Algorithm.
Let =$\mathcal{F}=\{x_{1},x_{2},\ldots,x_{d}\}$ denote the full set of features. The backward stepwise selection algorithm proceeds as follows:
  1. 1.
    $y=\beta_{0}+\sum_{i=1}^{d}\beta_{i}x_{i}+\varepsilon$ (3)
  2. 2.
    $t_{i}=\frac{\hat{\beta}_{i}}{\text{SE}(\hat{\beta}_{i})},$ (4)
  3. 3.
  4. 4.
  5. 5.
    Model Selection Criteria: At each iteration, evaluate the model using metrics such as the Akaike Information Criterion (AIC), Bayesian Information Criterion (BIC), Mallows’ $C_{p}$, Adjusted $R^{2}$, and Cross-Validation Error (CV Error). These metrics help balance model complexity and goodness of fit.
Evaluation Metrics for Our Wrapper Algorithm.
We employ multiple metrics to gauge not only how well each model fits the data, but also how efficiently it uses the available features. This multi-criteria evaluation helps us to verify a model that performs reliably, avoids overfitting, and remains computationally viable for energy-aware and thermal-critical environments.
    $\text{AIC}=2k-2\ln(L),$ (5)
    $\text{BIC}=k\ln(n)-2\ln(L),$ (6)
    $C_{p}=\frac{\text{RSS}}{\hat{\sigma}^{2}}-(n-2k),$ (7)
    Adjusted $\text{Adjusted }R^{2}=1-\left(\frac{(1-R^{2})(n-1)}{n-k-1}\right).$ (8)
By applying these evaluation metrics, we identify the optimal number of features that balance predictive accuracy and model simplicity. Figures 4(a) and 4(b) demonstrate that using fewer than 8 features suffices for accurate estimation of both average temperature and energy consumption, indicating that tracking only the most relevant predictors can improve energy efficiency and thermal behavior.
(a) Average temperature estimation
(b) Energy consumption estimation
(c) Determining the significance of features using Random Forest with respect to cross-validation error on Intel Core i7 12th Gen.
(d) Comparison of FCN models with and without feature selection and bootstrapping on Intel Core i7 12th Gen.
Figure 4: (Left) Backward stepwise selection for estimating energy consumption and average temperature. Retaining fewer than 8 predictors (features) yields accurate predictions in both cases. (Right) Feature importance analysis and comparison of FCN models with and without feature selection/bootstrapping. All experiments performed on Intel Core i7 12th Gen.

### v2

# Feature-Aware Task-to-Core Allocation in Embedded Multi-core Platforms via Statistical Learning

@@ Abstract
Optimizing task-to-core allocation can substantially reduce power consumption in multi-core platforms without degrading user experience. However, existing approaches overlook critical factors such as parallelism, compute intensity, and heterogeneous core types. In this paper, we introduce a statistical learning approach for feature selection that identifies the most influential features—such as core type, speed, temperature, and application-level parallelism or memory intensity—for accurate environment modeling and efficient energy minimization, a critical consideration for embedded systems. Our experiments, conducted with state-of-the-art Linux governors and thermal modeling techniques, show that correlation-aware task-to-core allocation lowers energy consumption by up to 10% and reduces core temperature by up to 5C compared to random core selection. Furthermore, our compressed, bootstrapped regression model improves thermal prediction accuracy by 6% while cutting model parameters by 16%, yielding an overall mean square error reduction of 61.6% relative to existing approaches. We provided results based on superscalar Intel Core i7 12th Gen processors with 14 cores, and validated our method across a diverse set of hardware platforms and effectively balanced performance, power, and thermal demands through statistical feature evaluation.
Keywords Energy Optimization $\cdot$ Embedded Systems $\cdot$ Heterogeneous Platforms $\cdot$ Statistical Feature Evaluation $\cdot$ Task-to-Core Allocation
@@ 1 Introduction
Task-to-core allocation is a pivotal technique for enhancing the performance and reliability of embedded systems, where workloads exhibit distinct thermal and power consumption behaviors depending on core assignments. In multi-core processors, allocations can involve high-performance cores, low-energy cores, and graphic processing units (GPUs), each with unique performance characteristics across various configurations. Improper allocation can lead to thermal overheating, triggering throttling and reducing chip reliability and lifespan Hosseinimotlagh and Kim (2019). Cooling costs for overheated chips are significant, approximately \$3 per watt of heat dissipation Skadron et al. (2003). While dynamic voltage and frequency scaling (DVFS) complements allocation by adjusting power dynamically—potentially reducing consumption by 75% without altering user experience Ratković et al. (2015)—our focus is on optimizing task-to-core allocation to manage thermal and energy constraints in embedded systems. Existing approaches often rely on historical workload and sensor data to predict future behavior Maity et al. (2022), but lack a systematic, statistically robust framework, which our hybrid methodology addresses with novelty and precision.
Embedded systems, foundational to automotive, telecommunications, and consumer electronics, must operate under strict power and thermal constraints while delivering high performance. Task-to-core allocation is crucial for managing these constraints in heterogeneous architectures with diverse core types (e.g., high-performance, low-power, GPUs). Feature selection and evaluation identify critical features and correlations, enabling optimal core assignments for thermal stability and energy efficiency Shekarisaz et al. (2021). For instance, in real-time automotive systems, allocation ensures compute-intensive tasks run on performance cores without overheating, while mobile devices assign power-hungry applications to high-performance cores and background tasks to low-power cores, extending battery life. Identifying features tied to power and thermal behavior—such as temperature, frequency, and core adjacency—establishes general rules for allocation, enhancing embedded system reliability. This paper introduces a hybrid statistical learning framework, integrating Random Forest (RF), backward stepwise selection, and correlation analysis, to optimize task-to-core allocation across platforms, prioritizing allocation over DVFS for embedded contexts.
Application performance depends on execution factors like core type, speed, temperature, and application characteristics—memory or compute intensiveness and parallelism level Pagani et al. (2018). Linux governors Brodowski et al. (2013) and related studies often focus on CPU utilization for DVFS policies, targeting latency, temperature, or energy Liu et al. (2021); Lin et al. (2023); Kim et al. (2021). However, for task-to-core allocation, characteristics like parallelism, memory/compute intensiveness, and branch counts are equally critical. Applications with frequent branch misses benefit from sequential execution on single cores, while parallel, compute-intensive tasks excel on GPUs or performance cores in heterogeneous platforms. Moreover, temperature correlations between cores indicate adjacency and overheating risks. Our approach transcends utilization-centric methods, employing a multi-stage statistical framework to capture a comprehensive feature set and their interdependencies, ensuring clarity and applicability in embedded systems.
Collecting real data post-execution for task-to-core allocation optimization is computationally expensive and imprecise due to hardware sampling delays. Instead, allocation decisions can leverage inference from trained environmental models that account for randomness and feature importance to mitigate overfitting Liu et al. (2021). An augmentation step further reduces sampling overhead, boosting efficiency. Complementary approaches include hierarchical multi-agent DVFS scheduling Pivezhandi et al. (2026b), zero-shot LLM-guided allocation Pivezhandi and Saifullah (2026b), flow-augmented few-shot RL for sensor data Pivezhandi and Saifullah (2026a), and graph-driven performance modeling Pivezhandi et al. (2026a). Our paper introduces the first statistical learning framework for embedded systems, uniquely integrating RF-based feature reduction, backward stepwise selection, and correlation-aware allocation with bootstrapping to create a robust, platform-agnostic model. Unlike prior heuristic or filter-based methods, which lack systematic feature evaluation or adaptability across heterogeneous platforms, our approach outperforms existing techniques—where they exist—achieving up to 10% energy savings and 61.6% lower mean squared error, as validated against SOTA baselines on diverse embedded hardware.
Little work applies statistical learning to feature selection for task-to-core allocation with the rigor we propose. Statistical learning predicts outcomes and reveals feature relationships. Prior efforts used extreme value theorem (EVT) for energy and execution time bounds in DVFS contexts Cazorla et al. (2019); Davis and Cucu-Grosjean (2019); Reghenzani et al. (2020); Pallister et al. (2017), while Liu et al. Liu et al. (2021) applied XGBoost for latency-related features, and Sasaki et al. Sasaki et al. (2007) used decision trees and correlation for energy per instruction. These lack the integrated, multi-method focus we introduce for allocation, synergizing RF, OLS, and correlation analysis for superior predictive power and generalizability in embedded systems.
Our framework shows that correlation-aware task-to-core allocation reduces chip temperature at intermediate utilization by leveraging thermally independent cores. Backward stepwise selection reveals a small feature subset suffices for accurate energy estimation, while RF extracts critical features with low overhead. Bootstrapping enhances dataset robustness, yielding a compact, efficient model. This hybrid methodology surpasses prior single-method approaches, validated across embedded platforms for practical impact.
The key contributions are:
@@
    This paper pioneers a hybrid statistical learning approach, integrating RF, backward stepwise selection, and correlation analysis, to evaluate feature importance for accurate environment modeling and compressed learning in task-to-core allocation for embedded systems.
@@
    We implement correlation-aware allocation, statistically reducing temperature and energy by assigning tasks to uncorrelated cores, validated across heterogeneous embedded platforms.
@@
    We employ bootstrapping to augment data and boost accuracy, supporting few-shot and meta-learning in resource-constrained embedded settings.
@@
    We rigorously test against state-of-the-art methods, achieving up to 10% energy reduction and 5°C temperature decrease. Our compressed, bootstrapped model cuts prediction error by 6%, reduces parameters by 16%, and improves mean squared error by 61.6% over SOTA. Data were extracted from Intel Core i7 8th and 12th Gen and Intel Xeon 2680 processors.
In the remainder, Section 5 surveys statistical learning and scheduling, Section 2 details motivation and challenges, and Section 3 describes the methodology. Section 4 evaluates three processors, followed by conclusions in Section 6.
@@ 2 Motivation and Challenges
Feature selection is a cornerstone of statistical learning, pivotal for optimizing task-to-core allocation in embedded multi-core systems. It addresses critical challenges—curse of dimensionality, large data management, and performance unpredictability—common to platforms like Intel Core i7 and NVIDIA Jetson TX2. To surpass prior filter-based approaches and establish novelty, we leverage a hybrid methodology integrating Random Forest (RF), backward stepwise selection, and correlation analysis, tailored to embedded constraints. This section clarifies these challenges in the context of task-to-core allocation, emphasizing generalizability across heterogeneous architectures and grounding our motivation in rigorous, platform-validated insights.
@@ 2.1 Curse of Dimensionality
Feature selection is essential to mitigate the curse of dimensionality, a challenge amplified in embedded multi-core systems. As feature count rises—encompassing frequency, temperature, and performance metrics like cache misses—the data space grows exponentially, leading to sparsity in high-dimensional spaces Li et al. (2017). This sparsity risks model overfitting, degrading predictive performance on unseen data. In heterogeneous embedded systems, the diversity of cores (performance, low-energy, GPUs) expands the state space, with each core type contributing unique thermal and performance characteristics. For instance, Intel Core i7 12th Gen’s hybrid P/E-cores and Jetson TX2’s big-LITTLE clusters introduce complex feature interactions. Efficient feature selection, as implemented via our RF-based reduction, reduces this complexity, enabling compact, generalizable models critical for real-time embedded applications—offering a novel advance over traditional dimensionality handling.
@@ 2.2 Large Data Management
Managing the growing volume of data is a pressing challenge for task-to-core allocation in embedded systems, where resource constraints demand efficiency. Allocation strategies may use static historical data or streaming inputs adapting to runtime conditions. With streaming data’s rise—common in automotive or mobile embedded platforms—memory management becomes critical, as unpredictable data volumes can overwhelm limited resources. Retaining unnecessary features, such as redundant performance counters, inflates storage and processing overhead, straining embedded systems. Our hybrid approach, combining RF for initial feature pruning and bootstrapping for data augmentation, optimizes processor models and allocation algorithms, reducing overhead while maintaining accuracy. This addresses large data challenges with clarity and applicability across static and dynamic embedded environments, surpassing simpler filter-based data handling.
@@ 2.3 Performance Unpredictability
Performance unpredictability in embedded multi-core processors stems from energy and thermal variability, driven by features like temperature, frequency, and core-specific performance data. Figure 2 illustrates this, showing energy consumption varying by an order of magnitude across identical frequency and core settings on Intel Core i7 8th Gen (4 cores), Intel Core i7 12th Gen (14 cores), and Intel Xeon 2680 v3 (12 cores) under OpenMP benchmarks. This variability underscores the need for feature selection to pinpoint significant predictors of energy and thermal behavior across platforms and configurations. Our methodology—using backward stepwise OLS for energy-critical features and correlation analysis for thermal independence—identifies robust predictors, enabling a global allocation policy. Unlike DVFS-centric prior works, we prioritize task-to-core allocation, validated rigorously with up to 10% energy savings and 5°C temperature reductions, ensuring practical impact in embedded systems.
Figure 1: Energy consumption variation across three different processors with identical frequency and core count settings: Intel Core i7 8th Gen (Corei78) with 4 cores, Intel Core i7 12th Gen (Corei712) with 14 cores, and Intel Xeon 2680 v3 (Xeon) with 12 cores. The results are shown for three different OpenMP benchmarks.
Figure 2: Determining the significance of features using Random Forest with respect to cross-validation error on Intel Core i7 12th Gen.
@@ 3 Design Methodology
Our objective is to develop a robust multistage methodology based on three feature selection approaches—embedded, wrapper-based, and filter-based—and show how their combined use informs an intelligent task-to-core allocation algorithm for embedded multi-core systems. We introduce a hybrid framework integrating Random Forest (RF), backward stepwise OLS, and Pearson correlation, surpassing traditional filter-based methods in efficiency and predictive power. We provide a structured workflow adaptable to diverse platforms like Intel Core i7 8th and 12th Gen, Intel Xeon 2680 v3, and NVIDIA Jetson TX2, while ensuring experimental rigor through validated outcomes (e.g., 10% energy savings, 5°C temperature reduction). This section details each stage, linking to embedded system optimization.
3.1 Data Collection and Environment Setup
We begin by profiling the target embedded hardware under diverse workloads to capture comprehensive data on energy consumption, temperature variations, and relevant performance metrics. The granularity of thermal sensors significantly varies across platforms. Intel Core i7 8th generation and Intel Xeon 2680 v3 processors provide homogeneous core architectures with individual per-core thermal sensors, enabling detailed per-core correlation analyses. In contrast, the Intel Core i7 12th generation features a heterogeneous architecture composed of performance (P-cores) and efficient (E-cores) cores, each type operating at different frequencies and performance characteristics. Similarly, the NVIDIA Jetson TX2 offers thermal readings aggregated at the cluster level, typically distinguishing between big and little core clusters due to its big-LITTLE architecture. Thus, our methodology accommodates per-core, heterogeneous core, and cluster-level sensor capabilities, specifically tailored to embedded computing environments, ensuring broad applicability.
Our data collection covers a wide range of configurations, including varying core counts (4, 12, or 14 cores), diverse CPU frequencies, and task prioritizations. Tasks include representative benchmarks such as Fibonacci (fib), matrix multiplication (Strassen), and sequence alignment (alignment), providing varied computational intensities and memory access patterns typical of embedded system workloads. This diverse dataset enables accurate modeling of core behaviors under different workloads and thermal conditions, essential for embedded system optimization.
Initially, the system setup involves configuring the frequency governor (typically schedutil) and initializing a temperature buffer capable of storing historical temperature data from cores (up to 10,000 entries). Additionally, a detailed task history is maintained, capturing execution parameters such as energy consumption, profiler metrics, and temperature data under varying task priorities and frequencies. This systematic variation ensures the dataset’s robustness and representativeness for embedded systems.
Energy monitoring employs built-in hardware interfaces such as /sys/class/powercap/intel-rapl/ on Intel processors or on-board power sensors for the Jetson TX2, recording total energy consumed or average power during task execution. Performance metrics collected include CPU frequency, utilization, memory bandwidth, and hardware performance counters like cache misses, gathered via utilities such as perf and cpufreq-info at consistent intervals.
The random core assignment strategy serves as a baseline to evaluate the effectiveness of our proposed correlation-aware core allocation strategy. In the random assignment approach, tasks run concurrently on randomly selected cores, typically half of the available cores excluding core 0, reserved for system tasks. Post-execution, we record metrics including temperature, energy consumption, and performance data. Execution is temporarily paused if core temperatures exceed predefined thresholds (e.g., depending on thermal throttling point of Intel Core i7 which is 100°C), allowing cores to cool down before resuming execution. This baseline facilitates comparative analysis against the correlation-aware allocation strategy.
The collected dataset serves as input for the subsequent embedded feature reduction process, employing Random Forest algorithms to identify minimal yet critical feature subsets. This refined feature set is then utilized for accurate environment modeling, guiding intelligent, temperature-correlation-driven task-to-core allocation strategies specifically designed for embedded systems. Ultimately, our structured, multi-stage methodology aims to optimize thermal management and energy efficiency across diverse embedded multi-core computing platforms.